[figure: 9.7 cm]

SPIN REORIENTATION TRANSITIONS IN COBALT

M.O. Steinitz and G.S. MacLeod^{a} aDepartment of Physics St. Francis Xavier University Antigonish, Nova Scotia B2G 2W5 Canada

We have studied the spin reorientation process in Cobalt using capacitance dilatometric measurements along the c- and a- axes of a single crystal. We find no evidence of critical behaviour at either of the temperatures at the beginning and end of this process. This suggests that the re-orientation process is a ”single-site” phenomenon, involving only changes in the local environment of each moment, rather than a co-operative phenomenon.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : oral

Corresponding author : Michael Steinitz

Address for correspondence : Dept. of Physics, St. Francis Xavier University, P.O. Box 5000, Antigonish, Nova Scotia B2G 2W5 Canada

Email address : msteinit@juliet.stfx.ca